OpenStudy (anonymous):

Find the point (x,0) that is equidistant from (6,1) and (-2, 5)

OpenStudy (anonymous):

\[d=\sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}}\] \[\sqrt{(6-x)^{2}+(1-0)^{2}}=\sqrt{(-2-x)^{2}+(5-0)^{2}}\]

OpenStudy (anonymous):

i think it is (0.5,0) OR (1/2,0)

OpenStudy (anonymous):

\[\sqrt{(6-x)^2+1} = \sqrt{(-2-x)^2+5^2} \] \[\sqrt{x^2-12x+37} = \sqrt{x^2+4x+29}\] squaring, 8=16x x=1/2